Problem:
In planning for a new item, a manufacturer assumes that the number of items produced x and the cost in dollars C of producing these items are related by a linear equation. Projections are that 100 items will cost $10,000 to produce and that 300 items will cost $22,000 to produce. Find the equation that relates C and x.
Hint: This is just like other linear equations. Pretend that C is y.
y = mx + b

Answer by stanbon(48546)

Projections are that 100 items will cost $10,000 to produce and that 300 items will cost $22,000 to produce. Find the equation that relates C and x.
--------
You have two points and you want the equation of the line.
Points: (100,10000) (300,22000)
slope = [22000-10000]/[300-100]=12000/200=60
--------
Form of the line C=mx+b; you need to find "b".
10000=60*100+b
b=4000
C=60x+4000
==============
Cheers,
Stan H.